OpenStudy (anonymous):

solve for x 12(x - 2) = 8x - 8 A. x = 4 B. x = 5 C. x = 2 D. x = 6

OpenStudy (anonymous):

hi welcome to openstudy

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ok so first take everything out of the parenthess

OpenStudy (anonymous):

so u would get 12x-24=8x-8 right?

OpenStudy (anonymous):

then combine like terms so subtract 8x from both sides

OpenStudy (anonymous):

so it would now be 4x-24=8

OpenStudy (shaik0124):

12x-24=8x-8 subtracr 8x on both sides 12x-24-8x=8x-8-8x 4x-24=-8 add 24 on both sides 4x-24+24=-8+24 4x=16 divide by4 on both sides x=4

OpenStudy (anonymous):

oops it would be 4x-24=-8

OpenStudy (anonymous):

dont give the answer @shaik0124 -_-

OpenStudy (anonymous):

thats not nice